4(2y+3)/5= 5(2-y)/2
Answer by Cromlix(4381)

4(2y+3)/5= 5(2-y)/2
8y + 12/5 = 10 - 5y/2
Denominator = 10
8y + 12/5 = 10 - 5y/2

2(8y + 12) = 5(10 - 5y)
----------------------
10
16y + 24 = 50 - 25y
--------------------
10
Collect like terms.
Change signs as you cross = sign
41y - 26
----------
10
Hope this helps.
:-)